# On the Convergence Analysis of Muon

### v1

@@ 4.1 Nonconvex case
Thus, $J_{t}$ is the sum of $r_{t}^{2}$ elements of $A_{t}\triangleq(U_{t}^{\top}\otimes V_{t}^{\top})H_{t}(U_{t}\otimes V_{t})$, where $A_{t}\in\mathbb{R}^{r_{t}^{2}\times r_{t}^{2}}$ can be viewed as a representation of $H_{t}$ under a certain congruence transformation depending on $U_{t}$ and $V_{t}$. In fact, if we assume $H_{t}=P_{t}\otimes Q_{t}$, where $P_{t}\in\mathbb{R}^{m\times m}$ and $Q_{t}\in\mathbb{R}^{n\times n}$, then $A_{t}=(U_{t}^{\top}P_{t}U_{t})\otimes(V_{t}^{\top}Q_{t}V_{t})$ and $J_{t}=\langle U_{t}^{\top}P_{t}U_{t},V_{t}^{\top}Q_{t}V_{t}\rangle\leq\|U_{t}^{\top}P_{t}U_{t}\|_{{\rm op}}\|V_{t}^{\top}Q_{t}V_{t}\|_{*}\leq\|P_{t}\|_{{\rm op}}\|Q_{t}\|_{*}$. Note that the complexity of GD for $\epsilon$-nuclear norm stationary point is $O(rL\Delta\epsilon^{-2})$, where $L$ is at least larger than $\max_{t}L_{t}$ with $L_{t}\triangleq\|H_{t}\|_{{\rm op}}=\max_{t}\|P_{t}\|_{{\rm op}}\|Q_{t}\|_{{\rm op}}$. Thus, when $H_{t}$ can be represented by $P_{t}\otimes Q_{t}$ and $Q_{t}$ (or $P_{t}$) are relatively low-rank such that $\|Q_{t}\|_{{\rm op}}\approx\|Q_{t}\|\ll r\|Q_{t}\|_{{\rm op}}$, $J=\frac{1}{T}\sum_{t}J_{t}\leq\frac{1}{T}\sum_{t}\|P_{t}\|_{{\rm op}}\|Q_{t}\|_{*}\ll r\max_{t}L_{t}\leq rL$, and the convergence rate of Muon can be better than GD’s.
However, various studies (Gur-Ari et al., 2018; Zhao et al., 2021, 2024; Cosson et al., 2023) have shown that the gradients during neural network optimization are also typically low rank. Thus, when we convert the stationary point of Frobenius norm to nuclear norm, the additional coefficient can be much smaller than $\sqrt{r}$ (as shown in Figure 2 (c,g)). Therefore, to compare the convergence rates of Muon and GD more precisely and more fairly, one should examine the ratios between the nuclear norm and Frobenius norm of their gradients as well as $J_{t}$ and $L_{t}$. For example, if, for $t=0,1,\ldots,T$,
we can claim Muon is better than GD for finding the stationary point of $f$.
We consider investigating and validating (4) in the optimization process of neural network training. For example, we consider the optimization process of a three-layer Multi-Layer Perceptron (MLP) for an image classification task on the MNIST dataset (LeCun et al., 1998). We optimize this neural network using GD and Muon (Algorithm 2) respectively, and record the loss, $J_{t}$, $L_{t}$, $\|\nabla f(W_{t})\|_{*}^{2}$ and $\|\nabla f(W_{t})\|_{\rm F}^{2}$ (with respect to one of the matrix parameters) during the optimization process. The results are shown in Figure 2. Detailed settings can be found in Appendix D. From Figure 2, we can observe that in the early stages of optimization, $J_{t}$ can actually be smaller than $L_{t}$ (b, f), while $\|\nabla f(W_{t})\|_{*}^{2}$ is significantly larger than $\|\nabla f(W_{t})\|_{\rm F}^{2}$ (d, g). As a result, Muon’s ratio ($\|\nabla f(W_{t})\|_{*}^{2}/J_{t}$) can be significantly larger than GD’s ($\|\nabla f(W_{t})\|_{\rm F}^{2}/L_{t}$) (d, h). Therefore, according to our theory in Equation 4, Muon can converge much faster than GD in the early phase, which is also corroborated in our experimental results (a, e). As the optimization approaches convergence, $J_{t}$ becomes comparable to or even larger than $L_{t}$ (b, f). However, throughout the entire optimization process of both GD and Muon, the ratio of Muon ($\|\nabla f(W_{t})\|_{*}^{2}/J_{t}$) remains consistently larger than that of GD ($\|\nabla f(W_{t})\|_{\rm F}^{2}/L_{t}$) (d, h), demonstrating the advantage of Muon over GD in optimizing neural network matrix parameters.
@@
Figure 2: Comparison of $J_{t}$, $L_{t}$, $\|\nabla f(W_{t})\|_{*}^{2}$ and $\|\nabla f(W_{t})\|_{\rm F}^{2}$ over the training process of GD and Muon (Algorithm 2). $f$ is defined as the cross-entropy loss of a MLP with three matrix parameters $W^{1}\in\mathbb{R}^{128\times 784},W^{2}\in\mathbb{R}^{64\times 128},W^{3}\in\mathbb{R}^{10\times 64}$ over a fixed subset of MNIST. We show the gradients and Hessians with respect to $W^{2}$ in this Figure. Detailed settings can be found in Appendix D.
4.2 Star-Convex case
(a) Function Value
(b) $\|W_{t}-W^{*}\|_{{\rm op}}$
(c) $D_{\rm F}^{2}L/(D_{{\rm op}}^{2}L_{*})$
Figure 3: Experiments on a quadratic function. Detailed settings can be found in Appendix D.
To further study the advantages of Muon, we consider the star convex case so that we can study the convergence of function value instead of the gradient norm.
Assumption 4.9 (Star convex).
$\langle\nabla f(W),W-W^{*}\rangle\geq f(W)-f^{*}.$
where $W^{*}\in\mathop{\rm argmin}_{W}f(W)$ is the optimal point.
Moreover, to ensure the convergence of Muon, we adopt the following mild assumption. Note that the same assumption is also used in Gupta et al. (2018); An et al. (2025) for their convergence analysis of their algorithms in similar settings.
Assumption 4.10.
For $W_{t},t=0,...,T$, generated by Algorithm 2, we assume $\|W_{t}-W^{*}\|_{{\rm op}}\leq D_{{\rm op}}$.
From Figure 3(b), we can observe that when we apply Muon on a quadratic function, $\|W_{t}-W^{*}\|$ is bounded, which aligns with 4.10.
With the spectral norm Lipschitz smoothness assumption, we have the following convergence theorem of Muon.
Theorem 4.11.
Under 4.9, 3.2 and 4.10, if we apply Algorithm 2 with $\eta_{t}=\eta$, then
Thus, to reach the precision $f(W_{T})-f^{*}\leq\epsilon$, we can take $\eta=O(\frac{\epsilon}{L_{*}D_{{\rm op}}})$, and the complexity is $O(L_{*}D_{{\rm op}}^{2}\epsilon^{-1}\text{log}\frac{\Delta}{\epsilon})$, where $\Delta=f(W_{0})-f^{*}$.
Moreover, if we apply Algorithm 2 with adaptive learning rates $\eta_{t}=\frac{\|\nabla f(W_{t})\|_{*}}{L_{*}}$, we have
The proof of Theorem 4.11 can be found in Appendix C. Moreover, similar to the last section, we also conduct convergence analysis of Muon under 3.1 and 4.7 in Appendix C.
Recall that, in the analysis of GD, to reach the precision $f(W_{T})-f^{*}\leq\epsilon$, the complexity is $O(LD_{\rm F}^{2}\epsilon^{-1})$, where $L$ is the smoothness constant under Frobenius norm and $D_{\rm F}=\|W_{0}-W^{*}\|_{\rm F}$ (Garrigos and Gower, 2023). Hence, to compare the convergence rate of Muon and GD, we need to compare $LD_{\rm F}^{2}$ with $L_{*}D_{{\rm op}}^{2}$. Similar to the discussions in the last section, when the Hessians of $f$ have a blockwise diagonal structure and are relatively “low rank”, i.e. if it satisfies 4.5 with $\Lambda$ such that $\|\Lambda\|_{{\rm op}}\approx\|\Lambda\|_{*}\ll r\|\Lambda\|_{{\rm op}}$, then $L_{*}D_{{\rm op}}^{2}\leq\|\Lambda\|_{*}D_{{\rm op}}^{2}\ll r\|\Lambda\|_{{\rm op}}D_{{\rm op}}^{2}$. In experiments, we usually find that $\|W_{t}-W^{*}\|_{{\rm op}}$ are decreasing as the $t$ increase (Figure 3(b)). Thus, we assume $D_{{\rm op}}\approx\|W_{0}-W^{*}\|_{{\rm op}}$. Though, the success of LoRA (Hu et al., 2022) in fine-tuning foundation models indicating a relatively low rank of $W_{0}-W^{*}$ is enough for fine-tuning, however, many evidences (Lialin et al., ; Jiang et al., 2024; Huang et al., 2025) demonstrate that a relatively high rank of $W_{0}-W^{*}$ is needed for pretraining and some complex fine-tuning tasks for large foundation models, low rank of $W_{0}-W^{*}$ can lead to degraded performance in those scenarios. Thus, we assume $W_{0}-W^{*}$ is relatively “high rank”. Then, we can expect $r\|W_{0}-W^{*}\|^{2}_{{\rm op}}\approx\|W_{0}-W^{*}\|^{2}_{\rm F}$ and $L_{*}D_{{\rm op}}^{2}\ll r\|\Lambda\|_{{\rm op}}D_{{\rm op}}^{2}\approx LD_{\rm F}^{2}$, which means Muon can have a better performance than GD.
In Figure 3, we conduct experiments comparing Muon and GD on quadratic function $f(W)=\text{tr}\left((W-W^{*})^{\top}Q(W-W^{*})\right)$, where $Q$ is a positive definite matrix with a poor condition number and is relatively “low rank”, i.e. $\|Q\|_{*}\approx\|Q\|_{{\rm op}}$. Note that the Hessian of $f$ is $I\otimes Q$, which satisfies 4.5 and is aligned with our pervious discussions. We randomly choose $W^{*}$, and set $W_{0}=0$. We can note that Muon outperforms GD in Figure 3 and most samples of ratio $D_{\rm F}^{2}L/(D_{{\rm op}}^{2}L_{*})$ are larger than 2, which validates the theoretical findings.
@@ 5 Conclusions and future directions
In this work, we established the convergence guarantees of Muon in both nonconvex and star convex settings. We provided detailed analysis and comparisons between Muon and GD, and established the conditions under which Muon can outperform GD. We showed that Muon can exploit the low-rank and approximately blockwise diagonal structure of Hessian matrices – a property commonly observed during neural network training. Our experiments on neural networks and quadratic functions supported and corroborated the theoretical findings. One of the limitations of our work is that our experiments are conducted on relatively small-scale neural networks, and one direction for future research is to investigate the structure of Hessian matrices in larger-scale neural networks. Moreover, studying the structures of Hessian matrices from theoretical perspectives and leveraging their structures to design new or better optimization methods are also interesting future directions.
@@
The Appendix is organized as follows. In Appendix A, we introduce some lemmas that will be utilized in the subsequent proofs and present the proof of Lemma 4.6. In Appendix B, we present the proofs of theorems in the nonconvex setting. In Appendix C, we present the proofs of theorems in the star convex setting. In Appendix D, we provide the detailed settings of our experiments. Moreover, in Appendix E, we provide an introduction to the Newton–Schulz iteration, which is commonly used in the practical implementation of Muon, and show how it approximates the orthogonalization process.
@@ Appendix A Lemmas
Suppose the spectral decomposition of $\Lambda$ is $\Lambda=U\Sigma U^{\top}$, where $U$ is an orthogonal matrix, $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},...,\sigma_{n})$, $\sigma_{1}\geq\sigma_{2}\geq...\geq\sigma_{n}$ are the singular values of $\Lambda$. Denote $AU=B=(b_{1}\,b_{2}\,...\,b_{n})$, where $b_{i}$ is the $i$-th column of $B$. Then, we have
@@
Thus, we proved Equation 5 and Equation 6. For Equation 7, we can prove it with the following inequalities
@@
For $t=0,1,...,T$, $M_{t}$ and $W_{t}$ generated by Algorithm 1, defining $C_{0}=\nabla f(W_{0})$, and when $t>0$, $C_{t}=\beta C_{t-1}+(1-\beta)\nabla f(W_{t})=(1-\beta)\sum_{i=1}^{t}\beta^{t-i}\nabla f(W_{i})+\beta^{t}\nabla f(W_{0})$, we have
@@
According to 3.3, we have following relationship about $G_{t}$ and $\nabla f(W_{t})$.

### v2

@@ 4.3.1 Comparison with GD
Thus, $J_{t}$ is the sum of $r_{t}^{2}$ elements of $A_{t}\triangleq(U_{t}^{\top}\otimes V_{t}^{\top})H_{t}(U_{t}\otimes V_{t})$, where $A_{t}\in\mathbb{R}^{r_{t}^{2}\times r_{t}^{2}}$ can be viewed as a representation of $H_{t}$ under a certain congruence transformation depending on $U_{t}$ and $V_{t}$.
To better illustrate the relationship between $J_{t}$ and $L$, we assume that $H_{t}$ can be expressed as a Kronecker product, i.e., $H_{t}=P_{t}\otimes Q_{t}$, where $P_{t}\in\mathbb{R}^{m\times m}$ and $Q_{t}\in\mathbb{R}^{n\times n}$. Let $\sigma_{p,t,1}\geq\sigma_{p,t,2}\geq\dots\geq\sigma_{p,t,m}$ denote the singular values of $P_{t}$, and $\sigma_{q,t,1}\geq\sigma_{q,t,2}\geq\dots\geq\sigma_{q,t,n}$ denote the singular values of $Q_{t}$. Then, we have $A_{t}=(U_{t}^{\top}P_{t}U_{t})\otimes(V_{t}^{\top}Q_{t}V_{t})$ and $J_{t}=\langle U_{t}^{\top}P_{t}U_{t},V_{t}^{\top}Q_{t}V_{t}\rangle\leq\sum_{i=1}^{r_{t}}\sigma_{p,t,i}\sigma_{q,t,i}$. Here, we use the Von Neumann’s trace inequality.
Note that in the nonconvex case, the complexity of GD for $\epsilon$-nuclear norm stationary point is $O(rL\Delta\epsilon^{-2})$, where $L$ is at least larger than $\max_{t}L_{t}$ with $L_{t}\triangleq\|H_{t}\|_{{\rm op}}=\|P_{t}\|_{{\rm op}}\|Q_{t}\|_{{\rm op}}$. Thus, when $H_{t}$ can be represented by $P_{t}\otimes Q_{t}$ and $Q_{t}$, $P_{t}$ are relatively low-rank such that $\sum_{i=1}^{r_{t}}\sigma_{p,t,i}\sigma_{q,t,i}\ll r\sigma_{p,t,1}\sigma_{q,t,1}$, then $J=\frac{1}{T}\sum_{t}J_{t}\leq\frac{1}{T}\sum_{t}\sum_{i=1}^{r_{t}}\sigma_{p,t,i}\sigma_{q,t,i}\ll\frac{1}{T}\sum_{t}rL_{t}\leq rL$, and the convergence rate of Muon can be better than GD’s. (Or one of $Q_{t}$, $P_{t}$ is relatively low-rank, i.e. if $Q_{t}$ is relatively low-rank such that $\|Q_{t}\|_{{\rm op}}\approx\|Q_{t}\|\ll r\|Q_{t}\|_{{\rm op}}$, we can also have $J=\frac{1}{T}\sum_{t}J_{t}\leq\frac{1}{T}\sum_{t}\|P_{t}\|_{{\rm op}}\|Q_{t}\|_{*}\ll r\max_{t}L_{t}\leq rL$.)
Similar to the previous discussions, in the star convex case, if $W_{0}-W^{*}$ is relatively “high rank” such that $rD_{\rm op}^{2}\approx\|W_{0}-W^{*}\|^{2}_{\rm F}$ and $Q_{t}$, $P_{t}$ are relatively low-rank such that $\tilde{J}\ll rL$, then we have $\tilde{J}D_{\rm op}^{2}\ll LD_{\rm F}^{2}$ and Muon can have a better performance than GD. For stochastic case and Algorithm 1, we have similar conclusions, see Section B.4 for more detailed discussions.
@@
Figure 3: Comparison of $J_{t}$, $L_{t}$, $\|\nabla f(W_{t})\|_{*}^{2}$ and $\|\nabla f(W_{t})\|_{\rm F}^{2}$ over the training process of GD and Muon (Algorithm 2) on a MLP model with three matrix parameters $W^{1}\in\mathbb{R}^{128\times 784},W^{2}\in\mathbb{R}^{64\times 128},W^{3}\in\mathbb{R}^{10\times 64}$. We show the gradients and Hessians with respect to $W^{2}$ in this Figure. Detailed settings can be found in Appendix E.
(a) Loss
(b) AdamW: $J_{t}$ and $L_{t}$
(c) AdamW: $\|\nabla f(W_{t})\|_{*}^{2}$ and $\|\nabla f(W_{t})\|_{\rm F}^{2}$
(d) AdamW: $\|\nabla f(W_{t})\|_{*}^{2}/J_{t}$ and $\|\nabla f(W_{t})\|_{\rm F}^{2}/L_{t}$
(e) Average ratio
(f) Muon: $J_{t}$ and $L_{t}$
(g) Muon: $\|\nabla f(W_{t})\|_{*}^{2}$ and $\|\nabla f(W_{t})\|_{\rm F}^{2}$
Figure 4: Comparison of $J_{t}$, $L_{t}$, $\|\nabla f(W_{t})\|_{*}^{2}$, $\|\nabla f(W_{t})\|_{\rm F}^{2}$ and the average ratio over the training process of a six-layer GPT-2 style model with AdamW and Muon. (e): The average ratio ($\|\nabla f(W_{t})\|_{*}^{2}L_{t}/(\|\nabla f(W_{t})\|_{\rm F}^{2}J_{t})$) of all matrix parameters optimized by Muon. In fact, not only does the average ratio of all matrix parameters satisfy Equation 7, but every matrix parameter optimized by Muon also satisfies Equation 7 at the steps tested in our experiments. Due to space limitations, in (b, c, d, f, g, h) we present records for the parameter “transformer.h.0.attn.c_attn.weight,” which is the attention matrix of the first attention layer. Detailed settings can be found in Appendix E.
[…]
$\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W))=\frac{1}{B}\sum_{i=1}^{B}\nabla^{2}l(Wh(x_{i}))\otimes(h(x_{i})h(x_{i})^{\top}).$ (6)
[…]
For multi sample case $B>1$ and more general nonconvex loss functions, theoretical analyses of $L$ and $J$ are difficult. Here, we conduct experiments to calculate the $L_{t}$ and $J_{t}$ alongside the optimization trajectory. Moreover, various studies (Gur-Ari et al., 2018; Zhao et al., 2021, 2024; Cosson et al., 2023) have shown that the gradients during neural network optimization are also typically low rank. Thus, when we convert the stationary point of Frobenius norm to nuclear norm, the additional coefficient can be much smaller than $\sqrt{r}$. Therefore, to compare the convergence rates of Muon and GD more precisely and more fairly, one should examine the ratios between the nuclear norm and Frobenius norm of their gradients as well as $J_{t}$ and $L_{t}$. For example, we can examine the following inequality:
If this inequality consistently holds alongside the optimization trajectory, Muon is expected to converge faster than GD.
We investigate and validate (7) during the optimization of a three-layer Multi-Layer Perceptron (MLP) trained on the MNIST dataset, as well as a six-layer GPT-2 style model trained on the Shakespeare character-level corpus. We optimize the MLP model using GD and Muon (Algorithm 2), and the GPT-2 model using AdamW and Muon. During training, we record the loss, $J_{t}$, $L_{t}$, $\|\nabla f(W_{t})\|_{*}^{2}$, $\|\nabla f(W_{t})\|_{\rm F}^{2}$ and calculate the ratio (7). The results are shown in Figure 3 and 4, with detailed settings provided in Appendix E. In both experiments (Figure 3 and Figure 4), we observe that Muon’s ratio ($\|\nabla f(W_{t})\|_{*}^{2}/J_{t}$) are consistently larger than GD’s ($\|\nabla f(W_{t})\|_{\rm F}^{2}/L_{t}$), demonstrating the advantage of Muon.
@@ 5 Conclusions and future directions
In this work, we have presented a comprehensive convergence rate analysis of Muon under various assumptions. We provided detailed comparisons between Muon and GD, and established the conditions under which Muon can outperform GD with different assumptions. We showed that Muon can exploit the low-rank structure of Hessian matrices, a property commonly observed during neural network training. Our experiments on neural networks and quadratic regressions supported and corroborated the theoretical findings. Promising directions for future research include investigating the structure of Hessian matrices from a theoretical perspective and leveraging these structural properties to develop new or more effective optimization methods.
@@
The Appendix is organized as follows. In Appendix A, we introduce some lemmas that will be utilized in the subsequent proofs, and we also give the proofs of Lemma 4.12 and Lemma A.3. In Appendix B, we present the proofs of theorems in the nonconvex setting. In Appendix C, we present the proofs of theorems in the star convex setting. In Appendix D, we present additional examples that illustrate the low-rank structure of real-world data matrices. In Appendix E, we provide the detailed settings of our experiments.
@@ Appendix A Lemmas
Suppose the spectral decomposition of $\Lambda$ is $\Lambda=U\Sigma U^{\top}$, where $U$ is an orthogonal matrix, $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{n})$ with $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{n}$ being the ordered singular values of $\Lambda$. Denote $AU=B=(b_{1}\,b_{2}\,\dots\,b_{n})$, where $b_{i}$ is the $i$-th column of $B$. Then, we have
@@
Thus, we have proved Equation 8 and Equation 9. For Equation 10, we can prove it with the following inequalities
@@
Lemma A.3.
Let $\Lambda\in\mathbb{R}^{n\times n}$ be positive definite. For a twice continuously differentiable function $f:\mathbb{R}^{m\times n}\to\mathbb{R}$, define $f_{v}(w)=f(W)$ with $w=\overline{\mathrm{vec}}(W)\in\mathbb{R}^{mn}$. Then the following are equivalent:
  1. 1.
    $f$ satisfies 4.11, i.e. for all $W,W^{\prime}\in\mathbb{R}^{m\times n}$,
  2. 2.
    For all $w\in\mathbb{R}^{mn}$,
    $I_{m}\otimes\Lambda\preceq\nabla^{2}f_{v}(w)\preceq I_{m}\otimes\Lambda.$
Proof.
Let $M:=I_{m}\otimes\Lambda$. For any $A\in\mathbb{R}^{m\times n}$, we have
$\text{tr}\left(A\Lambda A^{\top}\right)=\overline{\mathrm{vec}}(A)^{\top}(I_{m}\otimes\Lambda)\overline{\mathrm{vec}}(A)=\overline{\mathrm{vec}}(A)^{\top}M\overline{\mathrm{vec}}(A),$
and similarly
$\text{tr}\left(G\Lambda^{-1}G^{\top}\right)=\overline{\mathrm{vec}}(G)^{\top}M^{-1}\overline{\mathrm{vec}}(G).$
Moreover,
$\nabla f_{v}(w)=\overline{\mathrm{vec}}(\nabla f(W)).$
Hence the first condition is equivalent to
$\|\nabla f_{v}(w)-\nabla f_{v}(w^{\prime})\|_{M^{-1}}\leq\|w-w^{\prime}\|_{M},\qquad\forall w,w^{\prime}\in\mathbb{R}^{mn}.$
Here, for a vector $v\in\mathbb{R}^{d}$, its norm $\|v\|_{A}$ is defined as $\|v\|_{A}:=\sqrt{v^{\top}Av}=\|A^{1/2}v\|_{2}$, where $A\in\mathbb{R}^{d\times d}$ is positive definite.
We first prove $(2)\Rightarrow(1)$. For any $w,w^{\prime}\in\mathbb{R}^{mn}$, we have
$\nabla f_{v}(w)-\nabla f_{v}(w^{\prime})=\int_{0}^{1}\nabla^{2}f_{v}(w^{\prime}+t(w-w^{\prime})))\,(w-w^{\prime})\,dt.$
Therefore,
$M^{-1/2}(\nabla f_{v}(w)-\nabla f_{v}(w^{\prime}))=\int_{0}^{1}\bigl(M^{-1/2}\nabla^{2}f_{v}(w^{\prime}+t(w-w^{\prime}))M^{-1/2}\bigr)M^{1/2}(w-w^{\prime})\,dt.$
Under (2), we have
$I\preceq M^{-1/2}\nabla^{2}f_{v}(\cdot)M^{-1/2}\preceq I,$
so its operator norm is at most $1$. Thus
$\|M^{-1/2}(\nabla f_{v}(w)-\nabla f_{v}(w^{\prime}))\|_{2}=$ $\left\|\int_{0}^{1}\bigl(M^{-1/2}\nabla^{2}f_{v}(w^{\prime}+t(w-w^{\prime}))M^{-1/2}\bigr)M^{1/2}(w-w^{\prime})\,dt\right\|_{2}$
$\leq$ $\int_{0}^{1}\left\|M^{-1/2}\nabla^{2}f_{v}(w^{\prime}+t(w-w^{\prime}))M^{-1/2}\right\|_{\rm op}\|M^{1/2}(w-w^{\prime})\|_{2}\,dt$
$\leq$ $\|M^{1/2}(w-w^{\prime})\|_{2},$
which means
$\|\nabla f_{v}(w)-\nabla f_{v}(w^{\prime})\|_{M^{-1}}\leq\|w-w^{\prime}\|_{M}.$
Next we prove $(1)\Rightarrow(2)$. Fix any $w$ and any direction $u\in\mathbb{R}^{mn}$. Applying (1) to $w+tu$ and $w$ gives
$\|\nabla f_{v}(w+tu)-\nabla f_{v}(w)\|_{M^{-1}}\leq|t|\,\|u\|_{M}.$
Dividing by $|t|$ and letting $t\to 0$ yields
$\|\nabla^{2}f_{v}(w)u\|_{M^{-1}}\leq\|u\|_{M},\qquad\forall u.$
Let $v:=M^{1/2}u$, we have $u=M^{-1/2}v$ and
$\|M^{-1/2}\nabla^{2}f_{v}(w)M^{-1/2}v\|_{2}\leq\|v\|_{2},$
which implies
$\|M^{-1/2}\nabla^{2}f_{v}(w)M^{-1/2}\|_{{\rm op}}\leq 1.$
Since $\nabla^{2}f_{v}(w)$ is symmetric, this is equivalent to
$I\preceq M^{-1/2}\nabla^{2}f_{v}(w)M^{-1/2}\preceq I.$
Multiplying both sides by $M^{1/2}$ proves
$M\preceq\nabla^{2}f_{v}(w)\preceq M,$
i.e.,
$I_{m}\otimes\Lambda\preceq\nabla^{2}f_{v}(w)\preceq I_{m}\otimes\Lambda.$
This completes the proof. ∎
@@
For $t=0,1,\dots,T$, $M_{t}$ and $W_{t}$ generated by Algorithm 1, defining $C_{0}=\nabla f(W_{0})$, and when $t>0$, $C_{t}=\beta C_{t-1}+(1-\beta)\nabla f(W_{t})=(1-\beta)\sum_{i=1}^{t}\beta^{t-i}\nabla f(W_{i})+\beta^{t}\nabla f(W_{0})$, we have
@@
According to 3.3, we have the following relationship about $G_{t}$ and $\nabla f(W_{t})$.